# Ultra-Faint Milky Way Satellites Discovered in Carina, Phoenix, and Telescopium with DELVE Data Release 3

### v1

@@ 4.1 Properties from DECam Photometry
Since the three systems exhibit very different properties and have varying numbers and types of observable stars, we treat the isochrone parameters differently for each system. For the brightest system, Carina IV, which has the largest number of observable stars, we allowed all the isochrone parameters ($(m-M)_{0}$, $\tau$, $Z_{\rm phot}$) to vary, finding a distance modulus of $(m-M)_{0}=20.1_{-0.2}^{+0.1}$, age of $\tau=12.5_{-1.0}^{+0.3}$ and a metallicity of $Z_{\rm phot}=0.00013_{-0.00003}^{+0.00004}$. In the case of Phoenix III, freeing all parameters led to a bimodal posterior, with one peak at $Z_{\rm phot}=0.0001$, $\tau=12$ Gyr, and another more prominent peak at an unphysically young and metal-rich solution of $Z_{\rm phot}=0.0006$, $\tau=7$ Gyr. To eliminate this bimodality, we fixed the metallicity of the system to $Z_{\rm phot}=0.0001$, and we obtained a distance modulus of $(m-M)_{0}=20.3_{-0.1}^{+0.2}$, age of $\tau=11.8_{-0.4}^{+0.5}$ Gyr. For DELVE 7, we adopted the best-fit isochrone from deeper Magellan/IMACS photometry, with $\tau=10$ Gyr, $Z=0.0006$, and $(m-M)_{0}=18.1\pm 0.3$ (see Section 4.2).
@@
4.2 Properties of DELVE 7 from Magellan/IMACS Photometry
Figure 3: Follow-up imaging of DELVE 7 performed with Magellan/IMACS. (Left) $2\arcmin\times 2\arcmin$ $g$-band cutout image centered on DELVE 7, which can be seen as a clear overdensity of faint point-like sources. We overlay red ellipses demarcating $a_{h}$ and $2a_{h}$. (Right) Color–magnitude diagrams for point-like sources in the vicinity of DELVE 7. The left panel shows a circular region with $r<0\farcm 5$ while the right panel shows a equal-area background annulus with inner radius $r=3\arcmin$. Stars used for the isochrone fit are shown in blue; the best-fit isochrone (shown in red) was found to be $\tau=10$ Gyr, $Z=0.00061$, $(m-M)_{0}=18.1$. Although one bright candidate RGB member falls close to the isochrone, this star has a Gaia parallax of $\varpi=0.71\pm 0.15$ mas that places it well in the foreground of DELVE 7; we mark it here with an ‘x’ to emphasize its non-membership. The shaded regions show $\pm 0.2$ in color about the best-fit isochrone.
@@ 4.2 Properties of DELVE 7 from Magellan/IMACS Photometry
We performed PSF photometry on the stacked IMACS images using DAOPHOT and ALLFRAME (Stetson, 1987, 1994) following the procedure described in Mutlu-Pakdil et al. (2018). The photometry was calibrated to point sources in the DELVE DR3 catalog, including a color term, and was corrected for Galactic extinction following the procedure described in Section 2. The typical extinction at the position of DELVE 7 from Schlegel et al. (1998) is E(BV)=0.052E(B-V)=0.052, and we present the dereddened $g_{0}$ and $r_{0}$ magnitudes. In Figure 3 (right), we show the color–magnitude diagram of stars associated with DELVE 7 within $r<0.5\arcmin$ along with the color–magnitude diagram of stars in an equal-area background annulus. Figure 3 shows a sparse main sequence turn-off at $r\sim 21.5$ followed by a well-populated main sequence down to $g\sim 24.5$. Neither of these features are seen in the background region. We therefore confirm DELVE 7 as a bona fide new stellar system.
To obtain a more precise estimate of the properties of DELVE 7 from the IMACS imaging, we simultaneously fit the age, metallicity, and distance modulus of DELVE 7 using $\chi^{2}$ minimization over a restricted grid of PARSEC isochrone models spanning $10<\tau<13.5$ Gyr, $0.0001<Z<0.001$, $17<(m-M)_{0}<19$. For each set of parameters, we computed a 1D $\chi^{2}$ comparing the model colors to the observed sample of stars within $r<0\farcm 5$ from the system centroid ($\approx 2r_{h}$, as estimated from a preliminary ugali fit to the DECam data) with $g_{0}<24$, $(g-r)_{0}<1$. These cuts yield a pure selection of DELVE 7 members, shown in blue in Figure 3. The best-fit isochrone was found to be a $\tau=10$  Gyr, $Z=0.00061$, $(m-M)_{0}=18.09$ model. Given the extreme sparsity of the observed MSTO, however, the age and metallicity are largely unconstrained within our grid range, and thus these estimates are best considered nuisance parameters for the inference of the distance modulus. Based on the width of the main sequence in $g_{0}$ about the best-fit model, we visually estimated an approximate $\pm 0.3$ mag distance modulus uncertainty, i.e., $(m-M)_{0}=18.1\pm 0.3$. This translates to a physical distance of $D_{\odot}=42^{+6}_{-5}$ kpc.
5 Proper Motion and Metallicity Measurements of Carina IV
In this section, we present additional analysis that increase the confidence in the classification of Carina IV as an old, metal-poor, bound stellar systems rather than chance alignments of Milky Way stars. In Section 5.1, we use Gaia proper motion measurements to show that stars identified as Carina IV members exhibit consistent motions. While in Section 5.2, we use photometry from DECam grigri and CaHK filters to demonstrate that the members of Carina IV have low metallicities, consistent with old stellar populations.
@@ 5.1 Proper Motion
5.2 Photometric Metallicities
Carina IV was also imaged with the metallicity-sensitive, narrowband CaHK filter (N395) on DECam as part of the ongoing The Mapping the Ancient Galaxy in CaHK (MAGIC) survey (Chiti et al. in prep.). The N395 filter covers the Ca ii H & K lines and is similar to other metallicity sensitive filters (e.g., as used by the Pristine survey; Starkenburg et al., 2017). To derive metallicities, the survey follows the same approach as in Chiti et al. (2020, 2021), which involves forward-modelling fluxes using a grid of synthetic spectra generated with the Turbospectrum code (Alvarez & Plez, 1998; Plez, 2012) under the assumption of Local Thermodynamic Equilibrium (TSLTE). Further details on the TSLTE grid can be found in Barbosa et al. (2025). Analysis of the Sculptor dwarf spheroidal galaxy with MAGIC has determined that the metallicity precision achievable with the N395 filter in combination with DELVE broadband imaging can be as low as $\sigma_{\rm[Fe/H]}\sim 0.16$ (Barbosa et al., 2025), while the first high-resolution spectroscopic follow-up of six MAGIC candidates found a median $\Delta\rm{[Fe/H]}=-0.06$, confirming the high precision of the photometric estimates (Placco et al., 2025).
Carina IV was serendipitously observed by MAGIC with a single $1\times 720$ s image in the N395 filter, and an additional $2\times 720$ s dithers were obtained after Carina IV was identified as a dwarf galaxy candidate. For the N395 imaging, we are able to estimate metallicities close to the Gaia detection limit, achieving uncertainties of $\sigma_{\rm[Fe/H]}\sim 0.5$ at $g\sim 21$. To verify the identity of Carina IV, we examine the MAGIC metallicities for stars with high ugali membership probability. As shown in Figure 5, we find four stars with high ugali membership and low MAGIC metallicity (${\rm[Fe/H]}<-1.5$) similar to other ultra-faint dwarf galaxies (e.g., Simon, 2019; Fu et al., 2023). The mean metallicity determined by MAGIC for these four stars is $\overline{\rm[Fe/H]}=-2.0\pm 0.2$, which is on the high end of the ultra-faint dwarf galaxy population (Fu et al., 2023). However, the detection of low-metallicity stars confirms Carina IV as an old, metal-poor system and motivates spectroscopic follow up. The individual metallicity measurements for the high-ugali-membership stars are provided in Table 2 in Appendix A.
@@ 5.2 Photometric Metallicities
6 Discussion
Figure 6: (Left) Absolute $V$-band magnitude ($M_{V}$) versus azimuthally averaged physical half-light radius ($r_{1/2}$) for the three new Milky Way satellites and the population of known Local Group globular clusters and dwarf galaxies. (Right) The same diagram, but showing $M_{V}$ versus heliocentric distance ($D_{\odot}$). Spectroscopically confirmed dwarf galaxies and likely galaxy candidates are shown in purple circles, while Milky Way globular clusters are shown with green $\times$’ symbols. Ambiguous ultra-faint compact satellites are shown with orange pentagon symbols. The three systems discovered in this work are shown as colored stars with bars representing 68% posterior uncertainty intervals. Data on the Milky Way satellite population were compiled by Pace (2025), and we adopt the same classifications as Tan et al. (2025a).
@@ 6.1 Carina IV
Overall, Carina IV’s properties are consistent with confirmed satellite galaxies toward the faint end of the luminosity function. Specifically, Carina IV’s extended physical size ($r_{1/2}=40$ pc) is larger than almost all known star clusters. One notable feature of Carina IV is its proximity in projection to the Magellanic Clouds, lying at an angular distance of $\sim 11\fdg 6$ from the LMC, raising the possibility of an association, with 7 other systems already confirmed to be LMC satellites (Kallivayalil et al., 2018; Erkal & Belokurov, 2020; Patel et al., 2020; Correa Magnus & Vasiliev, 2022; Vasiliev, 2024; Pace et al., 2025).
@@
To test the potential association of Carina IV with the LMC, we carry out orbit modeling in a combined Milky Way + LMC potential following the methods of Erkal & Belokurov (2020). We use the McMillan (2017) potential for the Milky Way and for the LMC we assume a Hernquist profile with a total mass of $1.38\pm 0.255\times 10^{11}\,{\rm M_{\odot}}$ (Erkal et al., 2019) and a scale radius chosen to match the enclosed mass of $1.7\times 10^{10}\,{\rm M_{\odot}}$ at 8.7 kpc from van der Marel & Kallivayalil (2014). As the radial velocity of Carina IV is unknown, we vary the velocity between 200-200 to 800 km s-1 in steps of 5 km s-1. At each velocity we draw 100 samples from the Carina IV distance and proper motion errors, the LMC phase space and total mass, and the posterior for the Milky Way potential.
@@
We find a low probability of 1.8% for Carina IV to be associated with the LMC based on our orbit modeling. To evaluate this probability, we count the number of realizations in which Carina IV was energetically bound to the LMC at the end of orbit rewinding following Erkal & Belokurov (2020). Only the velocity range of $\sim 200-450~{\rm km~s^{-1}}$ has a multitude of bound samples. In this velocity range, the association probability varies between 2% and 9%, and the total probability is 7%. In the orbital models, we find an average relative distance of $49_{-9}^{+21}\,{\rm kpc}$ and relative velocity of $373_{-151}^{+153}$ km s-1 with respect to the LMC at Carina IV’s orbital pericenter with respect to the LMC, while the LMC’s escape velocity at $49~{\rm kpc}$ is $\sim 135$ km s-1. We note that the large tangential velocity errors ($\approx 200$ km s-1) make it difficult to accurately determine the orbit and any association with the LMC; however, these errors are expected to decrease by a factor of $\sim$2.8 in the upcoming Gaia DR4. Overall, Carina IV is more likely to be a Milky Way satellite behind the LMC system than to be part of the LMC system.
6.2 Phoenix III
@@
Due to its low luminosity, large heliocentric distance, and the absence of follow-up observations, Phoenix III is the least well characterized of the three newly discovered systems presented in this paper. In fact, at a heliocentric distance of $D_{\odot}=115^{+10}_{-7}$ kpc and an absolute $V$-band magnitude of $M_{V}=-1.2^{+0.7}_{-1.1}$, it is the faintest system discovered at a distance of >100{>}100 kpc to date (Pace, 2025). The dwarf galaxy candidates Sextans II ($D_{\odot}=126$ kpc; $M_{V}=-3.9$) and Virgo III ($D_{\odot}=151$ kpc; $M_{V}=-2.7$) were discovered at slightly farther distances in data from KiDS and HSC-SSP, but are both considerably brighter (Homma et al., 2024b; Gatto et al., 2024). The comparably faint Kim 2 ($D_{\odot}=105$ kpc; $M_{V}=-1.5$) is slightly closer and was discovered in early DECam data and confirmed by deep Gemini/GMOS imaging (Kim et al., 2015). Also comparable is the fainter, but slightly closer, Virgo I ($D_{\odot}=89$ kpc; $M_{V}=-0.7$), which was first identified in the deeper HSC-SSP survey (Homma et al. 2016; Crnojević et al. in prep.).
@@
The depth and uniformity of DES Y6 make it possible to measure half-a-dozen stars in the subgiant branch of Phoenix III and produce a relatively clean identification of a spatial overdensity. Visual inspection of color images from the Legacy Surveys Sky Viewer shows a collection of faint blue stars (Figure 1); however, the larger angular size of Phoenix III ($a_{h}=0\farcm 85$) makes it difficult to identify visually. The median physical size ($r_{1/2}=19$ pc), absolute magnitude ($M_{V}=-1.2$), and heliocentric distance ($D_{\odot}=115$ kpc) place Phoenix III in a region of parameter space that is occupied by both confirmed dwarf galaxies and star clusters, though more heavily populated by the former (Figure 6). However, Kim 2 ($D_{\odot}=105$ kpc, $M_{V}=-1.5$, $r_{1/2}=12$ pc) possesses similar properties and is suspected to be a star cluster (Kim et al., 2015). Moreover, given the large uncertainties in the size measurement of Phoenix III, the possibility that it is a star cluster cannot be ruled out.
@@
Despite residing in the DES footprint, Phoenix III was overlooked by previous searches of the DES data (Bechtol et al., 2015; Koposov et al., 2015b; Drlica-Wagner et al., 2015; Kim & Jerjen, 2015b; Drlica-Wagner et al., 2020; McNanna et al., 2024). This is due to the importance of the faint sub-giant branch stars at $g\sim 24$, which only became accessible with the most recent DES data release. In searches of the DES Y3 data, a hotspot was identified at the location of Phoenix III with a significance of $\sqrt{\rm TS}=4.9$ and ${\rm SIG}=5.5$, which did not exceed the threshold of that search ($\sqrt{\rm TS}>6$, SIG >> 6). It is likely that deeper imaging over this region of the sky by Rubin LSST will yield more systems similar to Phoenix III (Tsiane et al., 2025).
6.3 DELVE 7

### v2

@@ IV.1 Properties from DECam Photometry
Since the three systems exhibit very different properties and have varying numbers and types of observable stars, we treat the isochrone parameters differently for each system. For the brightest system, Carina IV, which has the largest number of observable stars, we allowed all the isochrone parameters ($(m-M)_{0}$, $\tau$, $Z_{\rm phot}$) to vary, finding a distance modulus of $(m-M)_{0}=20.1_{-0.2}^{+0.1}$, age of $\tau=12.5_{-1.0}^{+0.3}$ and a metallicity of $Z_{\rm phot}=0.00013_{-0.00003}^{+0.00004}$. In the case of Phoenix III, freeing all parameters led to a bimodal posterior, with one peak at $Z_{\rm phot}=0.0001$, $\tau=12$ Gyr, and another more prominent peak at $Z_{\rm phot}=0.0006$, $\tau=7$ Gyr. In the main analysis, we fixed the metallicity of the system to $Z_{\rm phot}=0.0001$, and we obtained a distance modulus of $(m-M)_{0}=20.3_{-0.1}^{+0.2}$, age of $\tau=11.8_{-0.4}^{+0.5}$ Gyr. We discuss the physical motivation for this choice in Section VI.2 and examine the alternative younger solution in more detail in Appendix B. For DELVE 7, we adopted the best-fit isochrone from deeper Magellan/IMACS photometry, with $\tau=10$ Gyr, $Z=0.0006$, and $(m-M)_{0}=18.1\pm 0.3$ (see Section IV.2).
@@
IV.2 Properties of DELVE 7 from Magellan/IMACS Photometry
Figure 3: Follow-up imaging of DELVE 7 performed with Magellan/IMACS. (Left) $2\arcmin\times 2\arcmin$ $g$-band cutout image centered on DELVE 7, which can be seen as a clear overdensity of faint point-like sources. We overlay red ellipses demarcating $r_{h}$ and $2r_{h}$. (Right) Color–magnitude diagrams for point-like sources in the vicinity of DELVE 7. The left panel shows a circular region with $r<0\farcm 5$ while the right panel shows a equal-area background annulus with inner radius $r=3\arcmin$. Stars used for the isochrone fit are shown in blue; the best-fit isochrone (shown in red) was found to be $\tau=10$ Gyr, $Z=0.00061$, $(m-M)_{0}=18.1$. The median Magellan photometric uncertainties for stars around DELVE 7, as a function of $g$-band magnitude, are indicated by red error bars. Although one bright candidate RGB member falls close to the isochrone, this star has a Gaia parallax of $\varpi=0.71\pm 0.15$ mas that places it well in the foreground of DELVE 7; we mark it here with an ‘x’ to emphasize its non-membership. The shaded regions show $\pm 0.2$ in color about the best-fit isochrone.
@@ IV.2 Properties of DELVE 7 from Magellan/IMACS Photometry
We performed PSF photometry on the stacked IMACS images using DAOPHOT and ALLFRAME (Stetson, 1987, 1994) following the procedure described in Mutlu-Pakdil et al. (2018). The photometry was calibrated to point sources in the DELVE DR3 catalog, including a color term, and was corrected for Galactic extinction following the procedure described in Section II. The typical extinction at the position of DELVE 7 from Schlegel et al. (1998) is E(BV)=0.052E(B-V)=0.052, and we present the dereddened $g_{0}$ and $r_{0}$ magnitudes. In Figure 3 (right), we show the color–magnitude diagram of stars associated with DELVE 7 within $r<0.5\arcmin$ along with the color–magnitude diagram of stars in an equal-area background annulus. Figure 3 shows a sparse main sequence turn-off at $g\sim 21.5$ followed by a well-populated main sequence down to $g\sim 24.5$. Neither of these features are seen in the background region. We therefore confirm DELVE 7 as a bona fide new stellar system.
To obtain a more precise estimate of the properties of DELVE 7 from the IMACS imaging, we simultaneously fit the age, metallicity, and distance modulus of DELVE 7 using $\chi^{2}$ minimization over a restricted grid of PARSEC isochrone models spanning $10<\tau<13.5$ Gyr, $0.0001<Z<0.001$, $17<(m-M)_{0}<19$. For each set of parameters, we computed a 1D $\chi^{2}$ comparing the model colors to the observed sample of stars within $r<0\farcm 5$ from the system centroid ($\approx 2.5\ r_{h}$, as estimated from the ugali fit to the DECam data) with $g_{0}<24$, $(g-r)_{0}<1$. These cuts yield a pure selection of DELVE 7 members, shown in blue in Figure 3. The best-fit isochrone was found to be a $\tau=10$  Gyr, $Z=0.00061$, $(m-M)_{0}=18.09$ model. Given the extreme sparsity of the observed MSTO, however, the age and metallicity are largely unconstrained within our grid range, and thus these estimates are best considered nuisance parameters for the inference of the distance modulus. Based on the width of the main sequence in $g_{0}$ about the best-fit model, we visually estimated an approximate $\pm 0.3$ mag distance modulus uncertainty, i.e., $(m-M)_{0}=18.1\pm 0.3$. This translates to a physical distance of $D_{\odot}=42^{+6}_{-5}$ kpc.
V Proper Motion and Metallicity Measurements of Carina IV
In this section, we present additional analysis that increase the confidence in the classification of Carina IV as an old, metal-poor, bound stellar systems rather than chance alignments of Milky Way stars. In Section V.1, we use Gaia proper motion measurements to show that stars identified as Carina IV members exhibit consistent motions. While in Section V.2, we use photometry from DECam grigri and CaHK filters to demonstrate that the members of Carina IV have low metallicities, consistent with old stellar populations.
@@ V.1 Proper Motion
For the other two candidates, the majority of the stars that possess high ugali membership probabilities ($p_{\rm ugali}>0.2$) are fainter than the Gaia detection limit of $G\sim 21$, preventing us from obtaining Gaia proper motion measurements (see Figure 2 and Appendix A for more details). However, future datasets, such as Vera C. Rubin Observatory’s Legacy Survey of Space and Time (LSST; Ivezić et al. 2019), may provide proper motion measurements for these fainter stars.
V.2 Photometric Metallicities
Carina IV was also imaged with the metallicity-sensitive, narrowband CaHK filter (N395) on DECam as part of the ongoing The Mapping the Ancient Galaxy in CaHK (MAGIC) survey (Chiti et al. in prep.). The N395 filter covers the Ca ii H & K lines and is similar to other metallicity sensitive filters (e.g., as used by the Pristine survey; Starkenburg et al., 2017). To derive metallicities, the survey follows the same approach as in Chiti et al. (2020, 2021), which involves forward-modelling fluxes using a grid of synthetic spectra generated with the Turbospectrum code (Alvarez and Plez, 1998; Plez, 2012) under the assumption of Local Thermodynamic Equilibrium (TSLTE). Further details on the TSLTE grid can be found in Barbosa et al. (2025). Analysis of the Sculptor dwarf spheroidal galaxy with MAGIC has determined that the metallicity precision achievable with the N395 filter in combination with DELVE broadband imaging can be as low as $\sigma_{\rm[Fe/H]}\sim 0.16$ (Barbosa et al., 2025), while the first high-resolution spectroscopic follow-up of six MAGIC candidates found a median $\Delta\rm{[Fe/H]}=-0.06$, confirming the high precision of the photometric estimates (Placco et al., 2025).
Carina IV was serendipitously observed by MAGIC with a single $1\times 720$ s image in the N395 filter, and an additional $2\times 720$ s dithers were obtained after Carina IV was identified as a dwarf galaxy candidate. For the N395 imaging, we are able to estimate metallicities close to the Gaia detection limit, achieving uncertainties of $\sigma_{\rm[Fe/H]}\sim 0.5$ at $g\sim 21$. To verify the identity of Carina IV, we examine the MAGIC metallicities for stars with high ugali membership probability ($p_{\rm ugali}>0.2$). As shown in Figure 5, we find four stars with high ugali membership and low MAGIC metallicity (${\rm[Fe/H]}<-1.5$) similar to other ultra-faint dwarf galaxies (e.g., Simon, 2019; Fu et al., 2023). The remaining five stars with MAGIC metallicity measurements are possibly contaminants from either the LMC halo or the Milky Way halo. The mean metallicity determined by MAGIC for these four stars is $\overline{\rm[Fe/H]}=-2.0\pm 0.2$, which is on the high end of the ultra-faint dwarf galaxy population (Fu et al., 2023). However, the detection of low-metallicity stars confirms Carina IV as an old, metal-poor system and motivates spectroscopic follow up. The individual metallicity measurements for the high-ugali-membership stars are provided in Table 2 in Appendix A.
@@ V.2 Photometric Metallicities
VI Discussion
Figure 6: (Left) Absolute $V$-band magnitude ($M_{V}$) versus azimuthally averaged physical half-light radius ($r_{1/2}$) for the three new Milky Way satellites and the population of known Local Group globular clusters and dwarf galaxies. (Right) The same diagram, but showing $M_{V}$ versus heliocentric distance ($D_{\odot}$). Spectroscopically confirmed dwarf galaxies and likely galaxy candidates are shown in purple circles, while Milky Way globular clusters are shown with green $\times$’ symbols. Ambiguous ultra-faint compact satellites are shown with orange pentagon symbols. The three systems discovered in this work are shown as colored stars with bars representing 68% posterior uncertainty intervals. Data on the Milky Way satellite population were compiled by Pace (2025), and we adopt the same classifications as Tan et al. (2025c).
@@ VI.1 Carina IV
Overall, Carina IV’s properties are consistent with confirmed satellite galaxies toward the faint end of the luminosity function. Specifically, Carina IV’s extended physical size ($r_{1/2}=40$ pc) is larger than almost all known star clusters. One notable feature of Carina IV is its proximity in projection to the Magellanic Clouds, lying at an angular distance of $\sim 11\fdg 6$ from the LMC, raising the possibility of an association, with 7 other systems already confirmed to be LMC satellites (Kallivayalil et al., 2018; Erkal and Belokurov, 2020; Patel et al., 2020; Correa Magnus and Vasiliev, 2022; Vasiliev, 2024; Pace et al., 2025).
We note in Fig. 1 and 2 an excess of sources just above the Carina IV MSTO at $(g-r)_{0}\sim 0.2$ and $g\sim 23$. Given Carina IV’s close proximity to the LMC, and the fact that these stars align well with the LMC main sequence, we interpret these sources as likely LMC halo main-sequence stars. This interpretation is further supported by the observation that when we apply an isochrone cut based on the LMC isochrone, we observe a background gradient with increasing density toward the direction of the LMC.
@@
To test the potential association of Carina IV with the LMC, we carry out orbit modeling in a combined Milky Way + LMC potential following the methods of Erkal and Belokurov (2020). We use the McMillan (2017) potential for the Milky Way and for the LMC we assume a Hernquist profile with a total mass of $1.38\pm 0.255\times 10^{11}\,{\rm M_{\odot}}$ (Erkal et al., 2019) and a scale radius chosen to match the enclosed mass of $1.7\times 10^{10}\,{\rm M_{\odot}}$ at 8.7 kpc from van der Marel and Kallivayalil (2014). As the radial velocity of Carina IV is unknown, we vary the velocity between 200-200 to 800 km s-1 in steps of 5 km s-1. At each velocity we draw 100 samples from the Carina IV distance and proper motion errors, the LMC phase space and total mass, and the posterior for the Milky Way potential.
@@
We find a low probability of 1.8% for Carina IV to be associated with the LMC based on our orbit modeling. To evaluate this probability, we count the number of realizations in which Carina IV was energetically bound to the LMC at the end of orbit rewinding following Erkal and Belokurov (2020). Only the velocity range of $\sim 200-450~{\rm km~s^{-1}}$ has a multitude of bound samples. In this velocity range, the association probability varies between 2% and 9%, and the total probability is 7%. In the orbital models, we find an average relative distance of $49_{-9}^{+21}\,{\rm kpc}$ and relative velocity of $373_{-151}^{+153}$ km s-1 with respect to the LMC at Carina IV’s orbital pericenter with respect to the LMC, while the LMC’s escape velocity at $49~{\rm kpc}$ is $\sim 135$ km s-1. We note that the large tangential velocity errors ($\approx 200$ km s-1) make it difficult to accurately determine the orbit and any association with the LMC; however, these errors are expected to decrease by a factor of $\sim$2.8 in the upcoming Gaia DR4. Overall, Carina IV is more likely to be a Milky Way satellite behind the LMC system than to be part of the LMC system.
VI.2 Phoenix III
@@
Due to its low luminosity, large heliocentric distance, and the absence of follow-up observations, Phoenix III is the least well characterized of the three newly discovered systems presented in this paper. In fact, at a heliocentric distance of $D_{\odot}=115^{+10}_{-7}$ kpc and an absolute $V$-band magnitude of $M_{V}=-1.2^{+0.7}_{-1.1}$, it is the faintest system discovered at a distance of >100{>}100 kpc to date (Pace, 2025). The dwarf galaxy candidates Sextans II ($D_{\odot}=126$ kpc; $M_{V}=-3.9$) and Virgo III ($D_{\odot}=151$ kpc; $M_{V}=-2.7$) were discovered at slightly farther distances in data from KiDS and HSC-SSP, but are both considerably brighter (Homma et al., 2024; Gatto et al., 2024). The comparably faint Kim 2 ($D_{\odot}=105$ kpc; $M_{V}=-1.5$) is slightly closer and was discovered in early DECam data and confirmed by deep Gemini/GMOS imaging (Kim et al., 2015). Also comparable is the fainter, but slightly closer, Virgo I ($D_{\odot}=89$ kpc; $M_{V}=-0.7$), which was first identified in the deeper HSC-SSP survey (Homma et al. 2016; Crnojević et al. in prep.).
@@
The depth and uniformity of DES Y6 make it possible to measure half-a-dozen stars in the subgiant branch of Phoenix III and produce a relatively clean identification of a spatial overdensity. Visual inspection of color images from the Legacy Surveys Sky Viewer shows a collection of faint blue stars (Figure 1); however, the larger angular size of Phoenix III ($a_{h}=0\farcm 85$) makes it difficult to identify visually.
As discussed in Section IV.1, the morphological fits for Phoenix III yield two distinct posterior peaks. One corresponds to an old stellar population with an age of $\sim$12 Gyr, while the other favors an intermediate-age population of $\sim$7 Gyr. The two solutions have similar structural parameters, including total magnitude, physical size, and heliocentric distance. We favor the old, metal-poor $\sim$12 Gyr solution, consistent with other old Milky Way ultra-faint satellites (Durbin et al., 2025). This preference is motivated by the fact that the metal-poor isochrone allows for the inclusion of the candidate blue horizontal branch (BHB) star, Gaia DR3 4959457079527646080 (see also Figure 7). The star has colors $g_{0}-r_{0}=-0.21\pm 0.01$ and $i_{0}-z_{0}=-0.05\pm 0.02$, which are consistent with a BHB classification from Belokurov and Koposov (2016), albeit with large photometric uncertainties due to its faintness. Moreover, using the method of Belokurov and Koposov (2016) to derive an estimate of $M_{g}$, we find that the BHB candidate has a distance modulus of $(m-M)_{0}=20.3$, which is consistent with that of Phoenix III. It is unlikely that a random Milky Way field BHB star is contaminating Phoenix III, as field BHBs are rare at such large Galactocentric distances (e.g., Fukushima et al., 2018; Thomas et al., 2018; Yu et al., 2024). However, since only a single BHB candidate is identified, we cannot rule out the possibility that it is instead a blue straggler or another type of contaminant. Therefore, for completeness, we also present the isochrone fits corresponding to the intermediate-age solution in Appendix B.
The median physical size ($r_{1/2}=19$ pc), absolute magnitude ($M_{V}=-1.2$), and heliocentric distance ($D_{\odot}=115$ kpc) place Phoenix III in a region of parameter space that is occupied by old, metal-poor dwarf galaxies and star clusters, though it is more heavily populated by the former (Figure 6). This motivated our classification of the system as a likely dwarf galaxy. However, Kim 2 ($D_{\odot}=105$ kpc, $M_{V}=-1.5$, $r_{1/2}=12$ pc) possesses similar properties and is suspected to be a star cluster (Kim et al., 2015). Moreover, given the large uncertainties in the size measurement of Phoenix III, the possibility that it is a star cluster cannot be ruled out.
@@
Despite residing in the DES footprint, Phoenix III was overlooked by previous searches of the DES data (Bechtol et al., 2015; Koposov et al., 2015; Drlica-Wagner et al., 2015; Kim and Jerjen, 2015b; Drlica-Wagner et al., 2020; McNanna et al., 2024). This is due to the importance of the faint sub-giant branch stars at $g\sim 24$, which only became accessible with the most recent DES data release. In searches of the DES Y3 data, a hotspot was identified at the location of Phoenix III with a significance of $\sqrt{\rm TS}=4.9$ and ${\rm SIG}=5.5$, which did not exceed the threshold of that search ($\sqrt{\rm TS}>6$, SIG >> 6). It is likely that deeper imaging over this region of the sky by Rubin LSST will yield more systems similar to Phoenix III (Tsiane et al., 2025).
VI.3 DELVE 7